OpenStudy (anonymous):

a father gave rs 500 to his 4 children. the eldest one got 2/5 of the money and the remaining amount of money was divided equally among the rest. find the amount of money the youngest child got.? i know the answer is 1/5 but i just want to know how to frame the solution urgent help

terenzreignz (terenzreignz):

@TayyabaRawjani Let's think of that 500 as some sort of pie (well, it kinda is) The eldest gets 2/5. The whole is 5/5. Once 2/5 is taken away, how much remains of the 5/5?

OpenStudy (anonymous):

3/5

OpenStudy (anonymous):

then each son got 1/5 portion

terenzreignz (terenzreignz):

That's right. And it's divided EQUALLY among the THREE remaining children (we're barring the eldest)

terenzreignz (terenzreignz):

So... that understood then?

OpenStudy (anonymous):

but i want to know how to frame this answr for a 6th grade student

terenzreignz (terenzreignz):

ahh. Well, it'd be easier to visualize if you immediately tell him/her what 2/5 of 500 is. It's 200, right?

terenzreignz (terenzreignz):

So... there'd be 300 remaining for the three other children. I guess it'd be simpler to work out how much is left for each if the remaining three get equal shares.

OpenStudy (anonymous):

yeah but its difficult for him to understand n make a written solution out of this

OpenStudy (anonymous):

i want a perfect and easy written solution for this step by step

terenzreignz (terenzreignz):

I don't know how his teacher usually solves things. Maybe a sample solution would help give us some sort of template :D

OpenStudy (anonymous):

i dont have a sample ryt now thats y asking for some help

terenzreignz (terenzreignz):

Are they on proportions?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

Well, a partitive proportion that adds up to 5 is 2:1:1:1 for four children.

OpenStudy (anonymous):

ok

terenzreignz (terenzreignz):

2 because the eldest gets 2/5 of the money.

terenzreignz (terenzreignz):

So, solving how they usually solve partitive proportions, we get: 2x + 1x + 1x + 1x = 500 where x is one share. Giving 5x = 500 x = 100 So, since the eldest takes two shares, eldest gets 200, while the rest have just one share, giving 100 each

OpenStudy (anonymous):

is it ok eldest got 200/500 =2/5 then the remaining amount was 300 which means each son got 100/500 = 1/5 so the youngest song got 1/5 of the money

terenzreignz (terenzreignz):

Helping a little brother out? :D

terenzreignz (terenzreignz):

It's a little vague :D